Circular No. 6470

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 R1 Carl W. Hergenrother, University of Arizona, reports his discovery of

a comet on films taken by Timothy B. Spahr, University of Florida, and himself at the University of Arizona’s site east of Tucson. The following observations are available:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.7.28437 0^{h}00^{m}12.^{s}45 +39^{o}20′45′′.8 14  Hergenrother
7.31296 0 00 00.49 +39 19 25.7  _{”}
8.13036 23 54 03.58 +38 38 45.7  _{”}
8.36182 23 52 21.78 +38 26 55.9 13.4 Offutt
8.36434 23 52 20.67 +38 26 47.8 13.2  _{”}
8.36612 23 52 19.87 +38 26 42.0 13.2  _{”}
8.36873 23 52 18.73 +38 26 33.6 13.0  _{”}
8.37292 23 52 16.91 +38 26 21.2 13.2  _{”}
8.37875 23 52 14.26 +38 26 03.4  Hergenrother
8.38361 23 52 12.17 +38 25 48.6  _{”}

C. W. Hergenrother and T. B. Spahr (Mount Bigelow). 0.41-m f/3 Schmidt. Comet well condensed, faint tail extending a few arcmin to the south. Their visual estimates on Sept. 8 suggest m_{1} ∼ 11–12.

W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien + CCD. Roughly circular coma ∼ 4′.25 across with 35′′ condensation. No tail on 60-second exposures.

SUPERNOVA 1986P IN NGC 5763 J. Mueller reports her discovery of an apparent supernova of mag ∼

17 on a IIIa-J plate taken with the 1.2-m Oschin Schmidt Telescope on 1986 July 2 by E. Helin and Mueller in the course of the second Palomar Sky Survey. The candidate is located 8′′ west and 6′′ south of the center of NGC 5763 (α = 14^{h}49^{m}.0, δ = +12^{o}29′, equinox 2000.0). A Palomar J plate taken on 1993 Mar. 3 by C. Brewer and Mueller shows nothing at the position; N. Reid also reports that an F plate taken on 1986 June 13 by W. Sargent and Mueller shows no object at the position, indicating the apparent supernova was < 3 weeks old when photographed. The image quality corresponds to nearby images indicating that it is very unlikely that the object is asteroidal in nature or a plate defect.

1996 September 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green